

********************************************************************
UARS WEEKLY STATUS REPORT - 05 APR 2002
********************************************************************
***********
SOLSTICE
***********
Nothing received.
***********
SUSIM
***********
SUSIM performed its daily mid-resolution and MgII solar scans during
periods of predicted UARS telemetry.
***********
ACRIM
***********
Nothing received.
***********
PEM
***********
Nothing received.
***********
HALOE
***********
HALOE returned to routine operations after power cycling during the peak of
the high beta period. HALOE was turned "off and on" six times as expected
during high beta periods and this time of year. NRT data was not always
available during this period which made it difficult to monitor instrument
performance. Use of the NRT's is even more important now with the loss of
"KCRT" real-time data as expected with the new "Cecil" processors that do
not have that capability. NRT data is necessary to insure the health of
HALOE, particular after a power on and high beta operations.
***********
MLS
***********
Nothing received.
***********
HRDI
***********
Nothing received.
***********
WINDII
***********
Nothing received.
***********
MPG
***********
The MPG attended the ECLIPSE Delta Operational
Readiness Review held on 4 April 2002.
The MPG continues to serve as the point-of-contact
for UARS instrument operations to support re-engineering
of the UARS Mission Operations Concept to develop
capabilities for the Science Traceability Mission.
Daily processing of instrument Daily Activity Plans
continues with no problems.
***********
CDHF
***********
Nothing received.
***********
FOT
***********
REPORTING PERIOD
This report covers 23 March 2002 (Orbit 57574, GMT Day 082) through 29
March 2002 (Orbit 57678, GMT Day 088).
SPACECRAFT OPERATION
The observatory is now in Forward Flight (Southern Hemisphere viewing)
and is performing nominally under two-battery operations. The
instrument operational changes were:
INSTRUMENT OPERATIONAL CHANGES
Instrument Time Orbit Comment
HALOE 082/2107 57587 Power OFF
HALOE 083/0643 57593 Power ON
HALOE 084/1519 57613 Power OFF
HALOE 085/1336 57627 Power ON
HALOE 086/1659 57644 Power OFF
HALOE 086/1922 57646 Power ON
HALOE 087/1335 57657 Power OFF
HALOE 087/1558 57659 Power ON
HALOE 088/1039 57670 Power OFF
HALOE 088/1334 57672 Power ON
The beta angle ranged between the angles of 50.39 degrees to 52.73
degrees for this report period. The beta angle is now increasing
towards a minimum angle of 0 degrees on 13 April 2002 (DOY 103).
Spacecraft battery 1, 2, and 3 performance monitoring continues.
Battery 1 remains OFF the charge relay. Battery 2 and Battery 3 are
currently maintaining greater than 25.44 V end-of-night (EON) load bus
voltage. Battery temperatures are stable with a temperature delta
between Battery 2 and Battery 3 of 0.94 to 1.25 degrees C. Additional
battery performance data for each battery on 23 March 2002 and 29
March 2002 are provided in Appendix D.
The clock error ranged from -11.1 to 11.2 msec during this report
period. Clock rate adjustments are listed in Appendix C.
All instrument and subsystem engineers were advised on the initial
excursion of ALERTS or Out-Of-Limits conditions detected by the FOT
for this report. A summary is listed below and the details of ALERT
and Out-Of-Limits occurrences are listed in Appendix B.
OUT-OF-LIMITS = 01
ALERTS = 00
The following real-time command sheets were executed for the listed
operational element(s) during this report period (see Appendix A).
FOT = 08
HALOE = 10
UARS Anomaly Reports initiated during this report period are listed in
Appendix E and summarized below.
FOT = 4
Ten (10) Anomaly Reports remained OPEN at the end of this report
period (see Appendix E).
TRANSPONDER FREQUENCIES:
Transponder A: 01/333 - 2287.496855
Transponder B: 02/043 - 2287.498879
GROUND SYSTEM OPERATION
Four (4) operational support problems occurred this period. Anomaly
Reports are included in Appendix E.
The data loss calculations are no longer being supplied by Data
Capture Facility (DCF).
OTHER SIGNIFICANT EVENTS
UARS SATELLITE OPERATIONS
At the Eclipse ORR on 21 March 2002 it was decided that full time
operations should begin on Eclipse. Full time operations continued
this week and only a few minor problems have been found. The FOT, in
conjunction with Raytheon software development, are in the process of
correcting these known problems. Other than these few issues, the
Eclipse system has been very stable and performed well in its first
few days of operations. A delta ORR is scheduled for 04 April 2002 to
discuss any outstanding issues and report FOT readiness for the loss
of MSOCC personnel on 05 April 2002.
MEETINGS
The FOT attended a UARS Eclipse Meeting on Thursday, March 28, 2002.
FUTURE SIGNIFICANT EVENTS
The next UARS Yaw slew is scheduled for 13 April 2002 (DOY 103). This
maneuver will be a Reverse to Forward Flight maneuver (Southern
Hemisphere viewing).
Attachments: Appendix A, Real-time Command Sheets
Appendix B, Out-Of-Limits & ALERT Occurrences
Appendix C, Clock & Solar Array Rate Corrections
Appendix D, UARS Battery Performance Data
Appendix E, UARS Anomaly Reports
APPENDIX A
REAL-TIME COMMAND SHEETS
Week ending 29 March 2002
ORBIT AOS NO. SUBSYS REASON
----- --- --- ------ ------
57587 082/2107 02-090 HALOE Power OFF (TBAT=33.9)
57593 083/0643 02-091 HALOE Power ON
57594 083/0820 02-092 OBC Clk Adj 0.00 to 20.6 msec/day
57612 084/1225 02-093 OBC Cleared TMONS (Eclipse)
57613 084/1519 02-094 HALOE Power OFF
57624 085/0903 02-095 OBC Clk Adj 20.6 to 0.00 msec/day
57627 085/1336 02-096 HALOE Power ON
57644 086/1659 02-097 ATC ATC LOCATION JUMP
57644 086/1659 02-098 OBC PROCCESSOR 18 CNTRL
57644 086/1659 02-099 OBC RTS 24 INH/ENA
57644 086/1659 02-100 HALOE Power OFF
57644 086/1659 02-101 CADH Power ON XMITTER A
57646 086/1922 02-102 HALOE Power ON
57655 087/1040 02-103 OBC/CLK Clk Adj 0.00 to 20.6 msec/day
57657 087/1335 02-104 HALOE Power OFF
57659 087/1558 02-105 HALOE Power ON
57670 088/1039 02-106 HALOE Power OFF
57672 088/1334 02-107 HALOE Power ON
APPENDIX B
OUT-OF-LIMITS & ALERTS OCCURRENCES
Week ending 29 March 2002
The following Out-Of-Limits occurrences were detected:
MNEMONIC STATE DESCRIPTION ORBITS
-------- ----- ----------- ------
HATBAT Y-HI Bolometer Area Temp 57578-9
The following ALERT occurrences were detected:
MNEMONIC DESCRIPTION ORBITS
-------- ----------- ------
-None this report period.
_____________________________________________________________________
APPENDIX C
CLOCK & SOLAR ARRAY RATE CORRECTIONS
Week ending 29 March 2002
CLOCK DRIFT RATE ADJUSTMENTS
ORBIT TIME CLOCK RATE CHANGE
----- ----- -----------------
57594 083/0820 Clk Adj 0.00 to 20.6 msec/day
57624 085/0903 Clk Adj 20.6 to 0.00 msec/day
57655 087/1040 Clk Adj 0.00 to 20.6 msec/day
APPENDIX D
UARS BATTERY PERFORMANCE DATA
Week ending 29 March 2002
23 March 2002 - GMT Day 082 Beta = 50.39 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ------- -------- ---------- ----- ------- ------
1 -33.6/-44.8 -3.16/-4.38 +0.00/-0.00 10.72 0.000 0.00
2 00.0/-100.8 +5.14/+4.48 +24.4/-14.8 25.60 1.040 19.92
3 +11.2/-67.2 +4.48/+3.17 +24.0/-14.4 25.44 1.040 19.46
29 March 2002 - GMT Day 088 Beta = 52.73 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ------- ---------- ---------- ----- ------- ------
1 -5.60/-22.4 -3.77/-4.38 +0.0/ -0.0 10.56 0.000 0.00
2 +11.2/-22.4 +4.81/+3.82 +24.0/-14.8 26.24 1.050 19.69
3 +11.2/-22.4 +3.82/+2.85 +23.2/-14.4 26.24 1.040 19.34
V/T LEVEL CHANGES
ORB# TIME LVL FR-TO
----- ---- ---------
-no changes V/T 5
APPENDIX E
UARS ANOMALY REPORTS
Week ending 29 March 2002
NEW ANOMALY REPORTS GENERATED
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
02-019 FOT 57579 082/0822 N/A AP Reboot
02-020 FOT 57591 083/0325 N/A AP 1 Crashed
02-021 FOT 57604 084/0022 N/A AP 1 Crashed
02-014 FOT 57624 085/0903 14170 Generic Late Acquisition
ANOMALY CLOSURE INFORMATION RECEIVED
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
02-015 FOT 57517 078/0510 24260 Late Acquisition
ANOMALY INVESTIGATION REPORTS GENERATED
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
-None this report period.
OPEN ANOMALY REPORTS
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
02-004 SOLSTICE 56883 036/0050 N/A Solstice in Wrong Mode
02-011 FOT 57466 074/2033 N/A AP-7 Disc Drive Failure
02-013 FOT 57467 074/2116 N/A AP-1 Hang-up
02-016 FOT 57523 078/1413 N/A AP-1 Halt/Gateway Fail
02-017 FOT 57553 080/1522 N/A Multiple TAC Problems
02-018 FOT 57556 080/1930 N/A AP-1 Crashed
02-019 FOT 57579 082/0822 N/A AP Reboot
02-020 FOT 57591 083/0325 N/A AP 1 Crashed
02-021 FOT 57604 084/0022 N/A AP 1 Crashed
02-014 FOT See List Mar 02 N/A Generic Late Acquisition